# Comparison of sEMG Encoding Accuracy Across Speech Modes Using Articulatory and Phoneme Features

### 20 Apr 2026 v1

@@ II-A Dataset and Preprocessing
We recruited twenty-four speech-typical participants, each producing $50$ distinct sentences drawn from the TIMIT corpus[8]. For every sentence, each participant repeated it three times in each speech mode, resulting in nine utterances per sentence per participant. The speech modes were defined as follows: aloud (normal voiced speech), mimed (overt articulation without phonation), and subvocal (attempted speech with an occluded vocal tract and no phonation).
@@ II-C Encoding Model
We denote by $\mathbf{w}\in\mathbb{R}^{F|\mathcal{L}|}$ the concatenated mTRF weights spanning all features and time lags for a given sEMG channel. We initially explored a wider lag window of [500,300][-500,300] ms and found no meaningful structure beyond $\pm 300$ ms; therefore, we used a final window of [300,300][-300,300] ms with 20 ms steps (matching 50 Hz sampling), i.e., $\mathcal{L}=\{-300,-280,\ldots,300\}$ ms. Let $X\in\mathbb{R}^{T\times F}$ denote the feature matrix at 50 Hz (either $A$, $P$, or [PA][P~A]), $y\in\mathbb{R}^{T}$ the envelope for one EMG channel, and $\mathcal{L}$ the set of lags. Let $t=1,\ldots,T$ index time samples at 50 Hz and $\tau\in\mathcal{L}$ denote a lag (in ms). We form a lagged design matrix
@@ III-A Encoding Across Modes and Electrodes
We performed cross-validated variance partitioning [10] using three models: articulatory-only ($A$), phoneme-only ($P$), and concatenation ($AP$). Denoting explained variances ($r^{2}$) by different features as $r^{2}_{A}$, $r^{2}_{P}$, and $r^{2}_{AP}$, we compute

### 27 Apr 2026 v2

@@ II-A Dataset and Preprocessing
We recruited twenty-four speech-typical participants, each producing 5050 distinct sentences drawn from the TIMIT corpus[8]. For every sentence, each participant repeated it three times in each speech mode, resulting in nine utterances per sentence per participant. The speech modes were defined as follows: aloud (normal voiced speech), mimed (overt articulation without phonation), and subvocal (attempted speech with an occluded vocal tract and no phonation).
@@ II-C Encoding Model
We denote by $\mathbf{w}\in\mathbb{R}^{F|\mathcal{L}|}$ the concatenated mTRF weights spanning all features and time lags for a given sEMG channel. We initially explored a wider lag window of [500,300][-500,300] ms and found no meaningful structure beyond $\pm 300$ ms; therefore, we used a final window of [300,300][-300,300] ms with 20 ms steps (matching 50 Hz sampling), i.e., $\mathcal{L}=\{-300,-280,\ldots,300\}$ ms. Let $X\in\mathbb{R}^{T\times F}$ denote the feature matrix at 50 Hz (either AA, $P$, or [PA][P~A]), $y\in\mathbb{R}^{T}$ the envelope for one EMG channel, and $\mathcal{L}$ the set of lags. Let $t=1,\ldots,T$ index time samples at 50 Hz and $\tau\in\mathcal{L}$ denote a lag (in ms). We form a lagged design matrix
@@ III-A Encoding Across Modes and Electrodes
We performed cross-validated variance partitioning [10] using three models: articulatory-only (AA), phoneme-only ($P$), and concatenation (APAP). Denoting explained variances ($r^{2}$) by different features as $r^{2}_{A}$, $r^{2}_{P}$, and $r^{2}_{AP}$, we compute